OpenStudy (anonymous):

I am having a hard time with 3 Algebra questions, heres one of them: - 1/2x + 2 = -x + 7 (I figured out one of the other questions, but I cant figure out these 2, the other one is below)

OpenStudy (anonymous):

About the other 3 questions, I figured out one of them, but the last one is: 4d − 12 = 12 − 4d solve for d. I really dont want answers, just help

OpenStudy (anonymous):



OpenStudy (anonymous):

nice flag >>.>>

OpenStudy (anonymous):

Haha, thanks @amber96

OpenStudy (anonymous):

No need to thank me for your gorgeous flag, I like a nice cup of bigotry when I wake up in the morning!

OpenStudy (unklerhaukus):

get all the x terms on one side of the equation

OpenStudy (anonymous):

>.>

OpenStudy (unklerhaukus):

\[- \frac12x + 2 = -x + 7\] add \(x\) to both sides take away \(2\) from both sides

OpenStudy (phi):

4d - 24 = -4d you want the terms with d on one side you have -24 on the left side, so "move" the 4d to the right side. you do this by adding -4d to both sides of the equation -4d + 4d -24= -4d-4d now simplify

OpenStudy (phi):

you should get -24= -8d now divide both sides by -8 to find d

OpenStudy (anonymous):

So, d is equal to 3